# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@
Abstract
We study the structure of wave functions in complex Chern–Simons theory on the complement of a hyperbolic knot, emphasizing the similarities with the topological string/spectral theory correspondence. We first conjecture a hidden integrality structure in the holomorphic blocks and show that this structure guarantees the cancellation of potential singularities in the full non-perturbative wave function at rational values of the coupling constant. We then develop various techniques to determine the wave function at such rational points. Finally, we illustrate our conjectures and obtain explicit results in the examples of the figure-eight and the three-twist knots. In the case of the figure-eight knot, we also perform a direct evaluation of the state integral in the rational case and observe that the resulting wave function has the features of the ground state for a quantum mirror curve.
@@ 1 Introduction
Chern–Simons (CS) theory with a complex gauge group has been an excellent laboratory for studying various aspects of quantum field theory (QFT) since it is essentially exactly solvable. In recent years, building on previous work by physicists and mathematicians, perturbative and non-perturbative methods have been introduced, making it possible to calculate various observables, and many beautiful and interesting results have been obtained. For example, for the complements of hyperbolic knots in the three-sphere, the wave function of the theory with gauge group $\text{SL}(2,{\mathbb{C}})$ has been defined rigorously ak , inspired in part by physics developments hikami07 ; dglz . This wave function satisfies, in addition, a difference equation, which can be determined by an appropriate quantization of the classical $A$-polynomial of the knot stavros-aj , as expected from physics arguments gukov .
In this paper, we will further study the wave functions for complements of hyperbolic knots in CS theory with gauge group $\text{SL}(2,{\mathbb{C}})$. As we will show, they share many structural similarities with the wave functions occurring in topological string theory and, more precisely, in the context of the so-called topological string/spectral theory (TS/ST) correspondence ghm ; cgm ; wzh ; mmrev and its open string version mz-wv ; mz-wv2 ; szabolcs ; fgrassi . Indeed, it has been found in both cases that the WKB expansion in $\hbar$ of the perturbative wave function of the topological string can be resummed into a $q$-series [^1] This is a resummation of a convergent series, so it does not involve the more sophisticated techniques of resummation of divergent series, like Borel resummation., where $q={\rm e}^{{\rm i}\hbar}$. However, this $q$-series displays singularities at all points of the form
@@
Since these singularities are not present in the non-perturbative definition of the wave function, they are an artifact of the resummed WKB expansion. This feature, in turn, requires the presence of a non-perturbative sector that cancels the singularities and leads to a finite answer for the full wave function. The cancellation of singularities at the values of $\hbar$ in Eq. (1.1) is a defining attribute of Faddeev’s quantum dilogarithm faddeev , which can be regarded as a simple example of a wave function in complex CS theory. A similar cancellation mechanism has appeared in a more complicated context in ABJM theory, where it is sometimes referred to as the HMO mechanism hmo2 , and is one of the facets of the non-perturbative proposal for topological string theory put forward in ghm ; cgm ; wzh . Moreover, the concrete realization of the cancellation of singularities in topological string theory is a consequence of the integrality structure of the topological string amplitudes. In the closed string case, one has the Gopakumar–Vafa integrality gv and its refinement hiv . In the open string case, as shown in kpamir ; mz-wv ; szabolcs , one needs an integrality property akin to the one found in ov ; lmv ; nawata . In this paper, we conjecture a similar integrality structure for the resummed WKB expansion of the wave functions of complex CS theory and support our claims with explicit evidence obtained in examples. This integrality structure, indeed, guarantees the cancellation of singularities. A similar integrality property has been found in ekholm , and it would be interesting to clarify its relationship with our results.
@@
Our integrality conjecture characterizes the so-called holomorphic blocks of the wave functions bdp , and the corresponding integer invariants can be calculated from these blocks when they are explicitly known. Analogously to the case of topological string theory, at the special values of $\hbar$ in Eq. (1.1), which we refer to as rational points, the generic formula for the exact wave function has an apparent singularity. However, after the cancellation mechanism takes place, one typically finds a relatively simpler expression. It is an interesting task to determine this expression for general rational numbers, as it was done for Faddeev’s quantum dilogarithm in garou-kas . In this paper, we develop various independent techniques to achieve this goal, following ideas proposed in the context of the open TS/ST correspondence butterfly ; hsx ; szabolcs . To begin with, one can use the underlying integrality structure to derive an explicit answer at rational values of $\hbar$ in terms of the newly introduced integer invariants. Second, one can start directly from the AJ equation for the wave function and specialize it to rational values, where one finds a quasi-periodic structure similar to the one appearing in spectral problems on lattices. Finally, one can directly evaluate the Andersen–Kashaev state integral at rational points using the techniques of garou-kas . We apply these three distinct methods and explicitly show that they lead to the same results in various examples.
@@
The paper is organized as follows. In Section 2, we review the necessary background notions on complex CS theory and the AJ conjecture. In Section 3, we present most of our results. In particular, we state the integrality conjecture for the WKB resummed wave function and show how it implies the cancellation of singularities at rational values of $\hbar$. We also provide two different techniques for evaluating the wave function at rational points. In Section 4, we illustrate these results by performing explicit computations in the concrete examples of the figure-eight and the three-twist knots. In Section 5, we evaluate directly the state integral in the rational case for the figure-eight knot. Finally, we conclude and list some open problems in Section 6. In the three Appendices, we provide additional details on the calculations we perform in Section 3 and recall some useful properties of Faddeev’s quantum dilogarithm.
2 Complex Chern–Simons theory and the $A$-polynomial
@@
In this section, we review the fundamental aspects of complex CS theory on a closed three-manifold and the construction of the classical and quantum $A$-polynomials. The physical understanding of the connection to A-polynomials was developed in gukov , and an excellent summary can be found in dimofte-rev . The AJ conjecture, which is at the basis of many of our computations, was proposed in stavros-aj . We pay special attention to the case of CS theory with $\text{SL}(2,{\mathbb{C}})$ gauge group on the complement of a hyperbolic knot in the three-sphere and introduce the two benchmark examples that we will consider later in this work, that is, the figure-eight and the three-twist knots.
2.1 Classical and quantum $A$-polynomials
@@
The classical action of CS theory with complex gauge group $G_{{\mathbb{C}}}$ can be written as witten91
$S=\frac{t}{8\pi}\int_{M}{\rm Tr}\left({\cal A}\wedge d{\cal A}+\frac{2}{3}{\cal A}\wedge{\cal A}\wedge{\cal A}\right)+\frac{t^{\prime}}{8\pi}\int_{M}{\rm Tr}\left(\bar{{\cal A}}\wedge d\bar{{\cal A}}+\frac{2}{3}\bar{{\cal A}}\wedge\bar{{\cal A}}\wedge\bar{{\cal A}}\right)\,,$ (2.2)
@@
where $M$ is the underlying closed three-manifold, whose boundary is the closed Riemann surface $\Sigma=\partial M$, ${\cal A}$ is the complex gauge field, which is a one-form on $M$ taking values in the Lie algebra $\mathfrak{g}_{{\mathbb{C}}}$, and $t,t^{\prime}$ are complex coefficients such that $t=k+s$ and $t^{\prime}=k-s$ where $k\in{\mathbb{Z}}$. For the theory to be unitary, $s$ must be real or purely imaginary, although we will not need this in our discussion. We also introduce the complex coupling constant ${\mathsf{b}}^{2}$ defined by
${\mathsf{b}}^{2}=\frac{t^{\prime}}{t}\,.$ (2.3)
The classical phase space of the theory is the moduli space of flat $G_{{\mathbb{C}}}$-connections on $\Sigma$ modulo gauge transformations, that is,
${\cal M}_{\rm{flat}}(G_{{\mathbb{C}}},\Sigma)={\rm Hom}\left(\pi_{1}(\Sigma),G_{{\mathbb{C}}}\right)/\sim\,,$ (2.4)
where $\pi_{1}$ denotes the fundamental group and $\sim$ is conjugation by elements of $G_{{\mathbb{C}}}$. This comes naturally equipped with a symplectic two-form
$\omega=\frac{t}{8\pi}\int_{\Sigma}{\rm Tr}\left(\delta{\cal A}\wedge\delta{\cal A}\right)+\frac{t^{\prime}}{8\pi}\int_{\Sigma}{\rm Tr}\left(\delta\bar{{\cal A}}\wedge\delta\bar{{\cal A}}\right)\,.$ (2.5)
@@
A classical solution on $M$ is a gauge equivalence class of flat $G_{{\mathbb{C}}}$-connections on $M$, which are gauge connections ${\cal A}$ satisfying the property
$d{\cal A}+{\cal A}\wedge{\cal A}=0\,,$ (2.6)
@@
and analogously for $\bar{{\cal A}}$. The moduli space of classical solutions on $M$ is a Lagrangian submanifold of the classical phase space. Namely, the submanifold
${\cal L}={\cal M}_{\rm{flat}}(G_{{\mathbb{C}}},M)\subset{\cal M}_{\rm{flat}}(G_{{\mathbb{C}}},\Sigma)$ (2.7)
is Lagrangian with respect to the symplectic structure $\omega$ in Eq. (2.5). In this paper, we will focus on a three-manifold $M$ obtained as the complement of a hyperbolic knot ${\cal K}$ in the three-sphere ${\mathbb{S}}^{3}$, that is,
@@
which has a single toral boundary $\Sigma={\mathbb{T}}^{2}$. Its moduli space of flat $G_{{\mathbb{C}}}$-connections is a complex Lagrangian submanifold ${\cal L}$ of the full phase space
${\cal M}_{\rm{flat}}(G_{{\mathbb{C}}},{\mathbb{T}}^{2})=\left({\mathbb{T}}_{{\mathbb{C}}}\times{\mathbb{T}}_{{\mathbb{C}}}\right)/{\cal W}\,,$ (2.9)
@@
$A_{{\cal K},i}(\bm{P},\bm{X})=0\,,\quad i=1,\dots,r\,,$ (2.10)
where the polynomials $A_{{\cal K},i}(\bm{P},\bm{X})$ have coefficients in ${\mathbb{R}}$.
Quantizing the classical phase space in Eq. (2.4) with its symplectic structure in Eq. (2.5) produces a Hilbert space ${\cal H}_{\Sigma}$. In the quantized theory, the path integral over the manifold $M$ leads to a vector $|M\rangle\in{\cal H}_{\Sigma}$, while the polynomials $A_{{\cal K},i}(\bm{P},\bm{X})$ in Eq. (2.10) are expected to produce quantum operators $\hat{A}_{{\cal K},i}(\hat{\bm{P}},\hat{\bm{X}},q)$, $i=1,\dots,r$, acting on ${\cal H}_{\Sigma}$, which annihilate the state $|M\rangle$. Here, the complex variables $\bm{P},\bm{X}$ are promoted to the operators $\hat{\bm{P}},\hat{\bm{X}}$, which satisfy the commutation relations
$\hat{P}_{i}\hat{X}_{j}=q^{\delta_{ij}}\hat{X}_{j}\hat{P}_{i}\,,\quad i,j=1,\dots,r\,,$ (2.11)
@@
is the complex coupling parameter playing the role of Planck’s constant. We will sometimes denote $\tau={\mathsf{b}}^{2}$. Note that the quantization of the theory depends on the level $k$ and, in this work, we will restrict ourselves to the case $k=1$. See, e.g.dimofte-levelk . The classical constraints in Eq. (2.10) become Schrödinger-like operator equations in the quantum theory. Namely,
$\hat{A}_{{\cal K},i}(\hat{\bm{P}},\hat{\bm{X}},q)\chi_{{\cal K}}(\hbar)=0\,,\quad i=1,\dots,r\,,$ (2.13)
where $\chi_{\cal K}(\hbar)$ is the partition function associated to the manifold $M$ in Eq. (2.8) defined by the complement of the knot ${\cal K}$. Following the conventions of ggm2 , we define the continuous complex variables $\bm{u}=(u_{1},\dots,u_{r})$ such that
$X_{i}={\rm e}^{2\pi{\mathsf{b}}u_{i}}\,,\quad i=1,\dots,r\,.$ (2.14)
@@
We stress that the quantization of the $A$-polynomials in Eq. (2.10) is not obtained simply by promoting $\bm{P},\bm{X}$ to their operator counterpart. On top of ordering issues, the resulting operators have a non-trivial dependence on $\hbar$ through $q$. It is fair to say that the spectral theory of these operators is not entirely understood, and a deeper grasp of this issue might clarify the corresponding non-trivial quantization problem. A concrete way of constructing the quantized $A$-polynomials is, for example, to use the original AJ conjecture of stavros-aj .
The wave function $\chi_{{\cal K}}(\bm{u};\hbar)$ can be calculated perturbatively in a saddle-point approximation where the saddle, or classical solution, of the EOM is described by the homomorphism
$\rho:\pi_{1}(M)\rightarrow G_{{\mathbb{C}}}\,.$ (2.16)
@@
Equivalently, the saddle points can be identified with the different classical solutions $\bm{P}^{(\alpha)}$ to Eq. (2.10) for fixed $\bm{X}$, which we denote with the additional discrete label $\alpha$. We recall that the branches of the polynomials in Eq. (2.10) come in conjugate pairs due to the symmetry of the theory under conjugation. Thus, each flat connection ${\cal A}^{(\rho)}$, labeled by $\rho\in{\rm Hom}\left(\pi_{1}(M),G_{{\mathbb{C}}}\right)$, has a conjugate flat connection ${\cal A}^{(\bar{\rho})}$, corresponding to the conjugate homomorphism $\bar{\rho}$. We will subsequently denote by $\bar{\alpha}$ the branch which is conjugate to $\alpha$. There is always an abelian branch $\alpha=\rm{abel}$ described by $P_{1}=\dots=P_{r}=1$, which is self-conjugate, and a geometric branch [^3] On the geometric branch, $\bm{u}$ can be thought of as parametrizing the quantum deformations of the complex hyperbolic structure of $M$. $\alpha=\rm{geom}$ containing the discrete faithful representation of $\pi_{1}(M)$ into $\text{PSL}(2,{\mathbb{C}})$, which has a distinct conjugate $\alpha=\rm{conj}$. In the rest of this work, we will only consider non-abelian branches. Therefore, we simplify the factor corresponding to the abelian classical solution from the polynomials in Eq. (2.10) and proceed to quantize the simplified form.
@@
When computed in the saddle-point approximation, that is, using the WKB method, the perturbative wave function, denoted by $\chi_{{\cal K}}^{(\alpha),\rm{WKB}}(\bm{u};\hbar)$ is given by an asymptotic power series in $\hbar$ of the form dglz
$\chi_{{\cal K}}^{(\alpha),\rm{WKB}}(\bm{u};\hbar)=\exp\left(\frac{1}{\hbar}S_{0}^{(\alpha)}(\bm{u})-\frac{\delta^{(\alpha)}}{2}\log\hbar+\sum_{n=0}^{\infty}S_{n+1}^{(\alpha)}(\bm{u})\hbar^{n}\right)\,,$ (2.17)
@@
where the leading coefficient $S_{0}^{(\alpha)}(\bm{u})$ is determined by the classical CS functional, the next-to-leading coefficient $\delta^{(\alpha)}$ is an integer, and the higher-order coefficients $S_{n}^{(\alpha)}(\bm{u})$, $n\geq 1$, are obtained, in principle, by summing systematically the contributions of Feynman diagrams with $n$-loops. Note that $\delta^{(\alpha)}=0$ for $\alpha\neq\rm{abel}$.
2.2 The case of $\text{SL}(2,{\mathbb{C}})$ gauge group

### v2

# On the Structure of Wave Functions in Complex Chern–Simons Theory

@@
[…]
On the Structure of Wave Functions in Complex Chern–Simons Theory
[…]
On the Structure of Wave Functions
in Complex Chern–Simons Theory
[…]
Marcos MARIÑO a and Claudia RELLA b
[…]
M. Mariño and C. Rella
[…]
a) Département de Physique Théorique and Section de Mathématiques,
[…]
Received February 10, 2025, in final form December 19, 2025; Published online January 06, 2026
[…]
We study the structure of wave functions in complex Chern–Simons theory on the complement of a hyperbolic knot, emphasizing the similarities with the topological string/spectral theory correspondence. We first conjecture a hidden integrality structure in the holomorphic blocks and show that this structure guarantees the cancellation of potential singularities in the full non-perturbative wave function at rational values of the coupling constant. We then develop various techniques to determine the wave function at such rational points. Finally, we illustrate our conjectures and obtain explicit results in the examples of the figure-eight and three-twist knots. In the case of the figure-eight knot, we also perform a direct evaluation of the state integral in the rational case and observe that the resulting wave function has the features of the ground state for a quantum mirror curve.
[…]
complex Chern–Simons theory; hyperbolic knots; $A$-polynomial; wave functions; holomorphic blocks; integrality structure
[…]
57K31; 81T45; 81Q05
@@ 1 Introduction
Chern–Simons (CS) theory with a complex gauge group has been an excellent laboratory for studying various aspects of quantum field theory (QFT) since it is essentially exactly solvable. In recent years, building on previous work by physicists and mathematicians, perturbative and non-perturbative methods have been introduced, making it possible to calculate various observables, and many beautiful and interesting results have been obtained. For example, for the complements of hyperbolic knots in the three-sphere, the wave function of the theory with gauge group $\mathrm{SL}(2,{\mathbb{C}})$ has been defined rigorously [2], inspired in part by physics developments [11, 27]. This wave function satisfies, in addition, a difference equation [4], which can be determined by an appropriate quantization of the classical $A$-polynomial of the knot [16], as expected from physics arguments [22].
In this paper, we will further study the wave functions for complements of hyperbolic knots in CS theory with gauge group $\mathrm{SL}(2,{\mathbb{C}})$. As we will show, they share many structural similarities with the wave functions occurring in topological string theory and, more precisely, in the context of the so-called topological string/spectral theory (TS/ST) correspondence [6, 21, 37, 42] and its open string version [15, 38, 39, 44]. Indeed, it has been found in both cases that the WKB expansion in $\hbar$ of the perturbative wave function of the topological string can be resummed into a $q$-series, [^1] This is a resummation of a convergent series, so it does not involve the more sophisticated summability techniques that apply to divergent series, like the Borel–Laplace resummation. where $q={\rm e}^{{\rm i}\hbar}$. However, this $q$-series displays singularities at all points of the form
@@
Since these singularities are not present in the non-perturbative definition of the wave function, they are an artifact of the resummed WKB expansion. This feature, in turn, requires the presence of a non-perturbative sector that cancels the singularities and leads to a finite answer for the full wave function. The cancellation of singularities at the values of $\hbar$ in equation (1.1) is a defining attribute of Faddeev’s quantum dilogarithm [13], which can be regarded as a simple example of a wave function in complex CS theory. A similar cancellation mechanism has appeared in a more complicated context in ABJM theory, where it is sometimes referred to as the HMO mechanism [24], and is one of the facets of the non-perturbative proposal for topological string theory put forward in [6, 21, 42]. Moreover, the concrete realization of the cancellation of singularities in topological string theory is a consequence of the integrality structure of the topological string amplitudes. In the closed string case, one has the Gopakumar–Vafa integrality [20] and its refinement [28]. In the open string case, as shown in [31, 38, 44], one needs an integrality property akin to the one found in [30, 36, 40]. In this paper, we conjecture an integrality structure for the resummed WKB expansion of the wave functions of complex CS theory and support our claims with explicit evidence obtained in examples. This integrality structure, indeed, guarantees the cancellation of singularities. A similar integrality property has been found in [12], and it would be interesting to clarify its relationship with our results.
@@
Our integrality conjecture characterizes the so-called holomorphic blocks of the wave functions [5], and the corresponding integer invariants can be calculated from these blocks when they are explicitly known. Analogously to the case of topological string theory, at the special values of $\hbar$ in equation (1.1), which we refer to as rational points, the generic formula for the exact wave function has an apparent singularity. However, after the cancellation mechanism takes place, one typically finds a relatively simpler expression. It is an interesting task to determine this expression for general rational numbers, as it was done for Faddeev’s quantum dilogarithm in [18]. In this paper, we develop various independent techniques to achieve this goal, following ideas proposed in the context of the open TS/ST correspondence [23, 25, 44]. To begin with, one can use the underlying integrality structure to derive an explicit answer at rational values of $\hbar$ in terms of the newly introduced integer invariants. Second, one can start directly from the AJ equation for the wave function and specialize it to rational values, where one finds a quasi-periodic structure similar to the one appearing in spectral problems on lattices. Finally, one can directly evaluate the Andersen–Kashaev state integral at rational points using the techniques of [18]. We apply these three distinct methods and explicitly show that they lead to the same results in examples.
@@
The paper is organized as follows. In Section 2, we review the necessary background notions on complex CS theory and the AJ conjecture. In Section 3, we present most of our results. In particular, we state the integrality conjecture for the WKB resummed wave function and show how it implies the cancellation of singularities at rational values of $\hbar$. We also provide two different techniques for evaluating the wave function at rational points. In Section 4, we illustrate these results by performing explicit computations in the examples of the figure-eight and three-twist knots. In Section 5, we evaluate directly the state integral in the rational case for the figure-eight knot. Finally, we conclude and list some open problems in Section 6. In the three appendices, we provide additional details on the calculations we perform in Section 3 and recall some useful properties of Faddeev’s quantum dilogarithm.
2 Complex Chern–Simons theory and the $\bm{A}$-polynomial
@@
In this section, we review the fundamental aspects of complex CS theory on a closed three-manifold and the construction of the classical and quantum $A$-polynomials of a hyperbolic knot. The physical understanding of the connection to $A$-polynomials was developed in [22], and an excellent summary can be found in [9]. The AJ conjecture, which is at the basis of many of our computations, was proposed in [16]. We pay special attention to the case of CS theory with $\mathrm{SL}(2,{\mathbb{C}})$ gauge group on the complement of a hyperbolic knot in the three-sphere and introduce the two benchmark examples that we will consider later in this work, that is, the figure-eight and three-twist knots.
2.1 Classical and quantum $\bm{A}$-polynomials
@@
The classical action of CS theory with complex gauge group $G_{{\mathbb{C}}}$ can be written as [43]
$S={{t\over 8\pi}}\int_{M}\operatorname{Tr}\left({\cal A}\wedge{\rm d}{\cal A}+{{2\over 3}}{\cal A}\wedge{\cal A}\wedge{\cal A}\right)+{{t^{\prime}\over 8\pi}}\int_{M}\operatorname{Tr}\left(\bar{{\cal A}}\wedge{\rm d}\bar{{\cal A}}+{{2\over 3}}\bar{{\cal A}}\wedge\bar{{\cal A}}\wedge\bar{{\cal A}}\right),$ (2.1)
@@
where $t$, $t^{\prime}$ are complex parameters, $M$ is the underlying closed three-manifold with boundary $\Sigma=\partial M$}, and $\cal A$} is the complex gauge field, which is a one-form on $M$ taking values in the Lie algebra $\mathfrak{g}_{{\mathbb{C}}}$. Here, $\bar{{\cal A}}$ denotes the complex conjugate of $\cal A$}. The coefficients $t$, $t^{\prime}$ can be written as $t=k+{\rm i}s$} and $t^{\prime}=k-{\rm i}s$. Here, quantization of the theory implies that the level $k$ is an integer, while unitarity requires $s$ to be either real or purely imaginary, although we will not use this in our discussion. We also introduce the complex coupling constant ${\mathsf{b}}^{2}$ defined by
${\mathsf{b}}^{2}={{t^{\prime}\over t}}.$ (2.2)
@@
A classical solution on $M$ is identified with a gauge equivalence class of flat $G_{{\mathbb{C}}}$-connections on $M$, which are gauge fields $\cal A$} satisfying the classical Euler–Lagrange equations
${\rm d}{\cal A}+{\cal A}\wedge{\cal A}={\rm d}\bar{{\cal A}}+\bar{{\cal A}}\wedge\bar{{\cal A}}=0.$ (2.3)
Because a flat $G_{{\mathbb{C}}}$-connection on $M$ is determined by a homomorphism
$\rho\colon\ \pi_{1}(M)\rightarrow G_{{\mathbb{C}}},$ (2.4)
the moduli space of classical solutions on $M$ is thus
${\cal M}_{{\rm flat}}(G_{{\mathbb{C}}},M)=\operatorname{Hom}(\pi_{1}(M),G_{{\mathbb{C}}})/{\sim},$ (2.5)
where $\pi_{1}$ denotes the fundamental group and $\sim$ is conjugation by elements of the gauge group $G_{{\mathbb{C}}}$. The classical phase space of the theory is instead given by the moduli space of flat $G_{{\mathbb{C}}}$-connections on the closed Riemann surface $\Sigma$ modulo gauge transformations, that is,
${\cal M}_{{\rm flat}}(G_{{\mathbb{C}}},\Sigma)=\operatorname{Hom}(\pi_{1}(\Sigma),G_{{\mathbb{C}}})/{\sim},$ (2.6)
which comes naturally equipped with the symplectic two-form
$\omega={{t\over 8\pi}}\int_{\Sigma}\operatorname{Tr}(\delta{\cal A}\wedge\delta{\cal A})+{{t^{\prime}\over 8\pi}}\int_{\Sigma}\operatorname{Tr}\bigl(\delta\bar{{\cal A}}\wedge\delta\bar{{\cal A}}\bigr).$ (2.7)
Recall that there is a natural map
$\iota\colon\ {\cal M}_{{\rm flat}}(G_{{\mathbb{C}}},M)\rightarrow{\cal M}_{{\rm flat}}(G_{{\mathbb{C}}},\Sigma)$ (2.8)
@@
induced by the inclusion $\pi_{1}(\Sigma)\hookrightarrow\pi_{1}(M)$. Indeed, the image of the moduli space of classical solutions in equation (2.5) under this map is a Lagrangian submanifold of the classical phase space. Namely,
${\cal L}=\iota({\cal M}_{{\rm flat}}(G_{{\mathbb{C}}},M))\subset{\cal M}_{{\rm flat}}(G_{{\mathbb{C}}},\Sigma)$ (2.9)
is Lagrangian with respect to the symplectic structure in equation (2.7).
In this paper, we will focus on a three-manifold $M$ obtained as the complement of a hyperbolic knot [^2] A three-manifold $M$ is hyperbolic if there is a discrete faithful representation of its fundamental group $\pi_{1}(M)$ into the group of orientation-preserving isometries of ${\mathbb{H}}^{3}$, that is, the group $\mathrm{Isom}^{+}\bigl({\mathbb{H}}^{3}\bigr)\cong\mathrm{PSL}(2,{\mathbb{C}})$. A hyperbolic knot is a knot whose complement in the three-sphere is hyperbolic. $\cal K$} in the three-sphere ${\mathbb{S}}^{3}$, that is,
@@
${\cal M}_{{\rm flat}}\bigl(G_{{\mathbb{C}}},{\mathbb{T}}^{2}\bigr)=({\mathbb{T}}_{{\mathbb{C}}}\times{\mathbb{T}}_{{\mathbb{C}}})/{\cal W},$ (2.11)
[…]
$A_{{\cal K},i}(\bm{P},\bm{X})=0,\qquad i=1,\dots,r,$ (2.12)
where the polynomials $A_{{\cal K},i}(\bm{P},\bm{X})$ have coefficients in $\mathbb{Z}$} [7].
Quantizing the classical phase space in equation (2.6) with its symplectic structure in equation (2.7) produces an infinite-dimensional Hilbert space ${\cal H}_{\Sigma}$, and the Feynman path integral over the manifold $M$ leads to a state $|M\rangle\in{\cal H}_{\Sigma}$. In the case of a hyperbolic knot complement, the polynomials $A_{{\cal K},i}(\bm{P},\bm{X})$ in equation (2.12) are expected to produce quantum operators $\hat{A}_{{\cal K},i}\bigl(\hat{\bm{P}},\hat{\bm{X}},q\bigr)$, $i=1,\dots,r$, acting on ${\cal H}_{\Sigma}$, which annihilate the state $|M\rangle$. Here, the complex variables $\bm{P}$, $\bm{X}$ are promoted to operators $\hat{\bm{P}}$, $\hat{\bm{X}}$ satisfying the commutation relations
$\hat{P}_{i}\hat{X}_{j}=q^{\delta_{ij}}\hat{X}_{j}\hat{P}_{i},\qquad i,j=1,\dots,r,$ (2.13)
@@
is the complex coupling parameter playing the role of Planck’s constant. We will sometimes denote $\tau={\mathsf{b}}^{2}$. Note that the quantization of the theory depends on the level $k$ through the parameter ${\mathsf{b}}^{2}$, and, in this work, we will restrict ourselves to the case $k=1$. See, e.g., [8].
The classical constraints in equation (2.12) become Schrödinger-like operator equations in the quantum theory. Namely,
$\hat{A}_{{\cal K},i}\bigl(\hat{\bm{P}},\hat{\bm{X}},q\bigr)\chi_{{\cal K}}(\hbar)=0,\qquad i=1,\dots,r,$ (2.15)
where $\chi_{{\cal K}}(\hbar)$ is the partition function associated with the manifold $M$ in equation (2.10). Following the conventions of [17], we define the continuous complex variables $\bm{u}=(u_{1},\dots,u_{r})$ such that
$X_{i}={\rm e}^{2\pi{\mathsf{b}}u_{i}},\qquad i=1,\dots,r.$ (2.16)
@@
We stress that the quantization of the $A$-polynomials in equation (2.12) is not obtained simply by promoting $\bm{P}$, $\bm{X}$ to their operator counterparts. On top of ordering issues, the resulting operators have a non-trivial dependence on $\hbar$ through $q$. It is fair to say that the spectral theory of these operators is not entirely understood, and a deeper grasp of this issue might clarify the corresponding non-trivial quantization problem. A concrete way of constructing the quantum $A$-polynomials is, for example, to use the original AJ conjecture of [16].
@@
The wave function $\chi_{{\cal K}}(\bm{u};\hbar)$ can be computed perturbatively in a saddle-point approximation where the saddle, or classical solution of the Euler–Lagrange equations, is described by the group homomorphism in equation (2.4). Equivalently, the saddle points can be identified with the different classical solutions $\bm{P}^{(\alpha)}$ to equation (2.12) for fixed $\bm{X}$, which we mark with the discrete label $\alpha$. We recall that the branches of the polynomials in equation (2.12) come in conjugate pairs due to the symmetry of the theory under conjugation. Thus, each flat connection ${\cal A}^{(\rho)}$, labeled by $\rho\in\operatorname{Hom}(\pi_{1}(M),G_{{\mathbb{C}}})$, has a conjugate flat connection ${\cal A}^{(\bar{\rho})}$, corresponding to the conjugate homomorphism $\bar{\rho}$. Consequently, we will denote by $\bar{\alpha}$ the branch conjugate to $\alpha$. Note that there is always an abelian branch $\alpha=\mathrm{abel}$ described by the equations $P_{1}=\dots=P_{r}=1$, which is self-conjugate, and a geometric branch [^4] On the geometric branch, $\bm{u}$ can be thought of as parametrizing the quantum deformations of the complex hyperbolic structure of $M$. $\alpha={\rm geom}$ containing the discrete faithful representation of $\pi_{1}(M)$ into $\mathrm{PSL}(2,{\mathbb{C}})$, which has a distinct conjugate $\alpha={\rm conj}$. In the rest of this work, we will only consider non-abelian branches. Therefore, we simplify the factor corresponding to the abelian classical solution from the polynomials in equation (2.12) and proceed to quantize the simplified form.
@@
When computed in the saddle-point approximation, that is, using the WKB method, around the classical solution labeled by $\alpha$, the perturbative wave function is denoted by $\chi_{{\cal K}}^{(\alpha),{\rm WKB}}(\bm{u};\hbar)$ and given by an asymptotic series in $\hbar$ of the form [11]
$\chi_{{\cal K}}^{(\alpha),{\rm WKB}}(\bm{u};\hbar)=\exp\left({{1\over\hbar}}S_{0}^{(\alpha)}(\bm{u})-{{\delta^{(\alpha)}\over 2}}\log\hbar+\sum_{n=0}^{\infty}S_{n+1}^{(\alpha)}(\bm{u})\hbar^{n}\right),$ (2.18)
@@
where the leading-order coefficient $S_{0}^{(\alpha)}(\bm{u})$ is determined by the classical CS functional in equation (2.1) up to integer multiples of $2\pi{\rm i}\bm{u}$, the next-to-leading-order coefficient $\delta^{(\alpha)}$ is an integer that vanishes when $\alpha\neq\mathrm{abel}$, and the higher-order coefficients $S_{n}^{(\alpha)}(\bm{u})$, $n\geq 1$, are obtained, in principle, by summing the contributions of $n$-loops Feynman diagrams systematically.
2.2 The case of $\bm{\mathrm{SL}(2,{\mathbb{C}})}$ gauge group